OpenStudy (anonymous):

Solve the system by the substitution method. Show your work. 2y - x = 5 x2 + y2 - 25 = 0 my answer: x: 2y - x = 5 2y - 5 = x so x = 2y - 5 -Plug this into 2nd equation: (2y - 5)² + y² - 25 = 0 -Use FOIL to solve the (2y - 5)² part: (2y - 5)(2y - 5) 4y² - 10y - 10y + 25 4y² - 20y + 25 So : 4y² - 20y + 25 + y² - 25 = 0 Which can be simplified to: 4y² + y² - 20y = 0 4y² + y² - 20y = 0 y(4y + y - 20) = 0 So, because of the 0 multiplication rule, y=0 x= -5 (plug in y=0 to original equations: 2y - x = 5 2(0) - x = 5, so x= -5) (-5,0)

OpenStudy (anonymous):

@hartnn my teacher said incomplete

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@master50777

OpenStudy (anonymous):

ill be back after lunch

hartnn (hartnn):

y(4y + y - 20) = 0 so, y = 0 or 4y+y-20 = 0

hartnn (hartnn):

5y -20 = 0 y =4

hartnn (hartnn):

when y =4, find x

OpenStudy (anonymous):

5y-20=0 5(4)-20=0 0=0 @hartnn

hartnn (hartnn):

x = 2y -5 when y=4 x = 8-5 =3

OpenStudy (anonymous):

is my work correct? and should i add to my previuos answer?

hartnn (hartnn):

yes, add that to your previous answer

hartnn (hartnn):

so your final 2 points are \(\Large (-5,0) \\ \Large (3,4)\)